OpenStudy (anonymous):

A physics student shoves a 0.50-kg block from the bottom of a frictionless 30.0° inclined plane. The student performs 4.0 J of work and the block slides a distance s along the incline before it stops. Determine the value of s.

OpenStudy (anonymous):

what is ur answer? i try to calulate but, it keep take wrong answer

OpenStudy (anonymous):

Work in relation to a force is defined as \[ W = \textbf F \cdot \textbf s = Fs \cos \theta \ ; \ \theta \ \textrm{is the angle between the force and the displacement} \\ \\ \ \ \ \ \ \ \textrm{when the work we want to look at is horizontal}\] If the force and the displacement aren't in completely the same direction, some of it doesn't get "translated" as work. In this case, the work that's being done is from gravity, which only acts in the y direction, so we use F s sin theta as a shorthand for F s cos (90 - theta) the angle from the vertical at the top of the ramp - at theta = 0, that is if the ramp were flat, Fs sin theta = 0 because there can be no work done by gravity on a flat surface.  ^_^

OpenStudy (anonymous):

wow awesome thanks

OpenStudy (anonymous):

is that what you meant?

OpenStudy (anonymous):

yep :)

OpenStudy (anonymous):

^_^

OpenStudy (anonymous):

I think essay way to understand when we use cos and sin. If the crate stay on the plat , then you use the work by W =F cos . S. If they fall down or go up we use sin. IS that correct, ?

OpenStudy (anonymous):